ㄈ 그 . ) The Brown Corporation is viewed as a possible takeover target y Cicron, Inc. Currentdy, Brown uses 20 percent debt in is cupital structure, but Cicron plans o increase the debt ratio to 25 percent if the scquistion is consummated The fer-tar cost of debt capital for Brown is estimated to be 8 percent, which holds constant under either apital structure. The cost of equty sfter th acquisition is eapected to be 22 percent. The current mar ket value of Browns oustanding debt is $75 million, ll f which will be assumed by Clcron Cicron intends to pay $225 million incash and common stock for lof Browns stock in addition to assuming all of Brown debt. Currently, the market price of Browns common stock is $200 million. Selected items from Browns financial data are as follows 2004 2005 2008 2007 THEREAFTER (MILLIONS) Net sales dministrative and slling expenses Depreciation Capital expenditures $260 $265 $280 $290 $300 25 15 17 8 23 30 22 18 18 20 22 2525 30 30 Inaddion, the cost of goods sold runs 50percent of sales and the marginal tar rate is Apercent. Compute the net present value of the acquisition Required: a. Compute NPV of the acquisition b. Make decision to acquire or not. Explain
Answer #1

In the given question, Cicron Inc (Acquirer) intends to acquire Brown Corporation (Target). To evaluate this proposal first we will find out the value of Brown Corporation by calculating the Net Present Value (NPV) by taking into account it's future cash flows after tax.

Given information

Particulars   Brown Corp   Cicron Inc (Post takeover)

Debt ratio 20% 25%

Cost of Debt (Kd ) 8% 8%

C ost of Equity (Ke) - 22%

Calculation of Weghted Average Cost of Capital (WACC)

WACC = (Cost of equity*Weight of equity)+(Cost of debt*Weight of debt)

= (Ke*We) + (Kd*Wd)

= (22*0.75) + (8*0.25)

= 18.5%

Calculation of Net Cash flows after tax (Amount in Million $)

Particulars 2004 2005 2006 2007 Thereafter
Sales 260 265 280 290 300
Less: Cost of Goods Sold (50%) (130) (132.5) (140) (145) (150)
Gross Profit 130 132.5 140 145 150
Less: Admin & Selling exp (25) (25) (25) (30) (30)
Less: Depreciation (15) (17) (18) (23) (30)
Profit Before Tax (PBT) 90 90.5 97 92 90
Less: Tax @34% (30.6) (30.77) (32.98) (31.28) (30.6)
Profit After Tax (PAT) 59.4 59.73 64.02 60.72 59.4
Add: Depreciation 15 17 18 23 30
Cash Flow After Tax 74.4 76.73 82.02 83.72 89.4
Less: Capital expenditure (22) (18) (18) (20) (22)
Net Cash Flow 52.4 58.73 64.02 63.72 67.4
Calculation of NPV (Amount in Million $)
Year 2004 2005 2006 2007 Terminal Period(Working Note)
Cashflow 52.4 58.73 64.02 63.72 364.32
Present Value factor@18.5% 0.844 0.712 0.601 0.507 0.507
Discounted Cashflow 44.23 41.82 38.48 32.31 184.71

Present value of future cash inflows = 44.23+41.82+38.48+32.31+184.71 = $ 341.55 mn

Present Cash outflow = Debts assumed + Consideration agreed to be paid

= $75 mn + $225 mn

= $ 300 mn

NPV = Present value of cash inflows - Cash outflow

= $ 341.55 mn -  $ 300 mn

= $ 41.55 mn

Decision to takeover

Since the NPV is positive, Cicron Inc. can takeover Brown Corporation.

Working Note

Since the cashflows after 2007 remain constant without any growth, it has to be brought down to 2007 period as follows.

Terminal period cashflow = Cashflow/ WACC

= 67.4/ 18.5%

= $ 364.32 mn